OpenStudy (keana):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 18 a: overlapping lines b: parallel lines c: intersecting lines

hero (hero):

Hint: Multiply first equation by 3

OpenStudy (anonymous):

check the slope!!! What can you say about the slope of these two equations!!

OpenStudy (keana):

i really don't understand it !

OpenStudy (anonymous):

Just do what Hero said

OpenStudy (keana):

i did but ...

hero (hero):

What happened?

OpenStudy (keana):

what am i looking for when i do that ?

OpenStudy (anonymous):

Post what you got then let us see it!!

hero (hero):

You're looking to see what happens when you multiply the first equation by three

OpenStudy (keana):

but am i solving for x or y or finding the slope and y intercept or what

hero (hero):

You're completely lost aren't you?

OpenStudy (anonymous):

3*(2x + y )= 3*6 6x + 3y = 18

OpenStudy (anonymous):

what do u get?

OpenStudy (keana):

3-y/2

hero (hero):

?

hero (hero):

What is that?

OpenStudy (keana):

y=6-2x & that

hero (hero):

Is that what you got after multiplying the first equation by 3?

OpenStudy (anonymous):

y=6-2x ok can you tell me the slope of this equation?

OpenStudy (keana):

m=-2, b=6

hero (hero):

2nd Hint: If you multiply the first equation by 3, you should get the second equation.

OpenStudy (anonymous):

Yes you got the slope right let's do it for the second one 6x + 3y = 18 3y = 18 - 6x y = (18 - 6x)/3 y = 3 - 2x remember y=mx+b can you tell me the slope of the second equation

OpenStudy (keana):

ok hold on

OpenStudy (anonymous):

You are almost there!! so close!

OpenStudy (keana):

m=-2 b=6 for 6x + 3y = 18

OpenStudy (anonymous):

Yes so m and b are the same for both equations!!!

OpenStudy (anonymous):

If m are the same that means that the equations are parallel or overlapping but if both b and m are the same that means that both equations are overlapping!!!

OpenStudy (anonymous):

This basically means that both equations are the same !!! Do you understand now?

OpenStudy (keana):

yes

hero (hero):

I'm saddened that you didn't understand my hint

OpenStudy (anonymous):

Hey everyone does it their own way @Hero